OpenStudy (anonymous):

What was the hiker's average speed for the entire hike? A. 0.187 km/hr B. 2.6 km/h C. 5.3 km/h D. 4 km/h

OpenStudy (anonymous):

km/hr\[\frac{24}{\frac{9}{2}}=\frac{16}{3}=5.33333 \]